SOUL COM Night #2 UT - 01 Oct 2020

OSA: D.G. Huerta

ISp: G. Bechetti

Mtn. Manager: B. Bingham

ISA. B. Rothberg (secret bunker)

AO: G. Taylor, J. Power


NightPlan:

  • OSCO
  • Pal10 or M92
  • GRB host galaxy
  • Performance Testing

Problems:

  1. Trouble closing the loops 1st half of the night.
  2. Offsets from large sky moves caused skip frames. They were all from the same regions of the sky. Hypothesis: if the offs are close enough that active optics is running, there might be a delay when it hands back off to AO on the return
  3. Had a brownout around 11:30PM local time. Knocked a lot of things out (mirror, air compressors, etc)


Night Log:

UT 00:18 - starting up LUCI GUIs and init' all

UT 00:19 - taking test darks

L1 0001-0002 2.51 sec

L1 0003 - checking fieldstop alignment - J. Power using robs to test that RTD works

L1 0004 - confirmatory image

UT 01:03 - DG.Hueta going through OSCO steps (G. Taylor started up the WFS already) before opening

UT 01:14 - opening enclosure
OSCO

UT 01:19 - opticsieve got put in FPU, should be the N30 fieldstop (screenshot in the instructions shows opticsieve)

-OSCO google doc screenshot has been fixed

UT 01:34 - sending OSCO preset to HD 230519

-star not found, need to do pointing and collimation first

UT 01:49 - GS for OSCO star was way out of focus, going back to another star to do a collimation check

UT 02:00 - took a while to get collimated, not sure why. guider/WFS show 0.6" before closing the loop

--oops, lower floods in chamber were on while trying to close the loop. they are now off

AO Info:

Lower floods on during AO acquisition so checkflux may be impacted. El 75, sent magnitude 7.5, OCAM magnitude 8.7 , bin 1, freq 1700, average pupil intensity 41, photons/subapp 50.3
L1 0005-0024
OSCO Part Deux

UT 02:27 - DIMM 1.1"

moving to BD +38 3564

AO Info: El 83, sent magnitude 7.51, OCAM magnitude 8.7 , bin 1, freq 1700, average pupil intensity 40, photons/subapp 49.0

UT 02:20 - DIMM 0.91"

L1 0025-0044
Pal 10

UT 02:30 - going to Pal 10 with PA=36 (last night's was PA = 306 to avoid a re-rotator error)

UT 02:31 - guider/WFS show 0.7"/0.66"

AO Info: El 74, sent magnitude 12.06, OCAM magnitude 13.4 , bin 2, freq 898, average pupil intensity 35, photons/subapp 7.0

L1 0045- fieldstop badly shifted, fixing now

L1 0046-0047 - verification of N30 fieldstop

L1 0048-0071

UT 02:39 - L1 0048 - repeat manually set up for first image

DIMM 0.85"

UT 03:08 - DIMM at zenith is 0.91
M52

UT 03:14 - moving to M52

UT 03:17 - guider/WFS 0.73"/0.71"

AO Info: El 50, sent magnitude 12.35, OCAM magnitude 13.8 , bin 2, freq 823, average pupil intensity 26, photons/subapp 5.1, DIMM 0.95"

L1 0072- first exposure jumped a bit vertically - just 1/26 frames - integrated looks OK

UT 03:24 - DIMM @zenith 1.16" EL 47 DIMM=1.47" (current EL is 51)

UT 03:29 - guiding on off-field guider/WFS show 0.835"/0.73"

L1 0076 - Safe skipping - this is a return from a large offset - DO NOT USE

-In auto mode now, we can't recover without sending a new preset!!!

-have to hit resume first and then send 0,0 Absolute DETXY offset

L1 0072-0075 ok to use

UT 03:33 - resending preset - will pick up after large offset, but this will re-take an extra 0,0 dither position

-the offs in this scrpt all are in the same area, so if this is directional, we will know

AO Info: El 52, sent magnitude 12.35, OCAM magnitude 13.7 , bin 2, freq 827, average pupil intensity 27, photons/subapp 5.6,

UT 03:43 - skipping to where we left off

DIMM 0.9"

UT 03:48 - guider/WFS (on off but still working) 0.81"/0.65"

small safe skip on this return, 1 skip frame detected otherwise OK

FWHMs ~0.067"

UT 03:54 - safe skips again on return from off (frame L1 0083)

third time seeing this. all offsets are same region (+200 X -200 Y)

FWHM H-band 0.058"

UT 04:04 - guding on offset, guider/WFS show 1"/-0.9"

DIMM 0.9"

-no safeskip or wanings occur from offset

UT 04:12 - bad safe skip frame L1 0093 (H-band) - do not use

UT 04:15 - recovered and continuing

UT 04:23 - offset to sky (active guiding) guider/WFS show 0.73:/0.62"

L1 0077-0100

-only covered K and H-band, skipped J-band

Could the safeskips be caused because the off fields still have active guiding - transition from active guiding back to AO the problem? (J. Power)

GRB 200826A

PSF star first

UT 04:27 - sending preset to PSF star (El 60)

do see some thin cirrus illuminated by Moon off to the East

AO Info: El 61, sent magnitude 15.27, OCAM magnitude 16.3 , bin 3, freq 310, average pupil intensity 18, photons/subapp 3.4, DIMM 1.17

UT 04:36 - DIMM 1.76" @zenith

DIMM upt to 2"

UT 04:40 - waiting a couple of minutes to see if this is a bubble or bad seeing. If bad will go back to M52 and collect data in poor conditions

UT 04:45 - bubble seems to be subsiding...

UT 04:47 - cancelling AO preset and resedning LUCI preset - went back down to 1.1-1.2"

AO info: El 61, sent magnitude 15.27, OCAM magnitude 16.5 , bin 4, freq 319, average pupil intensity 18, photons/subapp 3.4, DIMM 1.57

UT 04:48 - went back up to 1.85"

UT 04:49 - holding....
DIMM.jpg

L1 0101 - garbage frame

UT 05:01 - sending 0,0 Absolute DETXY offset

UT 05:05 - seeing back down to 1" - continuin script (20 min downtime)

FWHM PSF star L0103 = 0.39" - might not be usable - need to revisit after science field

L1 0102-0109

UT 05:17 - sending preset to GRB host galaxy field

- pointing check needed

UT 05:21 - resending script

Guider/WFS shows 0.65"/0.63"

UT 05:27 - OCAM issues?

UT 05:34 - had to cancel preset - G. Taylor having some issues (bandwidth?)

-OCAM looks to be down, pupil check would not work, restarting the WFS arbitratot

J. Power noticed earlier that OCAM appeared to havae gibberish, this was likely the first sign of an issue

-BCU2K in funky state

UT 05:37 - need to do wstop wstart

downtime - G. Taylor warming and cooling OCAM

-putting in blind filter and turning flexure off since until everything is back up again...

UT 06:02 - fixing N30 fieldstop

L1 0110 - check

L1 0111-0113 - verification

OK

UT 06:06 - sending preset to science field again

guider/WFS shows 0.89"/0.7"

UT 06:19 - back up, powered cycled BCU again

UT 06:20 - Shell RIP'd

-WFS arbitrator thinks loop closed, AdSec thinks we're in failure

UT 06:33 - elevation access halted - explains why vanished off guider

Why does it show ARGOS lockouts on my GUI???

UT 06:35 - air compressors off

J. Power cancelled preset

Looks like we had a brownout - faillures everywhere

UT 06:37 - putting in blinds and flexure off

SX mirror panic

UT 06:37 - running init all on both LUCIs - should clear ARGOS lockout (spurious) - cleared

- will check fieldstop later

UT 06:45 - D.G.Huerta resetting air compressors

UT 06:58 - enclosure running, compressors running, working on primary mirror now

UT 07:02 - should probably check AO system - G.Taylor says everything looks OK

UT 07:07 - D.G. Huerta working on additional systems to get us back online

Never got to GRB Host :(

-will roll over to tomorrow night at 10pm Local time (currently says TBD so there's an open slot)

Performance Tests

UT 07:20 - checking pointing

Next target AO121, PA=180 R=8.39

-testing Bin 1

UT 07:39 - closing loop

AO Info: AO121 sent Rmag 8.39mag, OCAM 9.8, bin1 freq 1700, average pupil intensity 47, 17.4 photons/subapp

test exposure FeII/clear, 256x256 42k peak, Rotator = 309 degrees
UT 08:07 - sending 5" DETX offset to get off frame

UT 08:10 - running modal machine tests

UT 08:44 - adSec arbitrator went offline, J, Power checking on it - got back without losing the loop

UT 08:51 - DIMM is 0.95"

UT 08:53 - resending preset

-recentered star on LUCI sub-frame

-J. Power having some trouble with optical gain

UT 09:13 - 60Hz vibration seen

UT 09:15 - moving +5" DETX for sky image

First offset done with MGM on

-lost connection from arbitrator to LUCI for exposures

UT 09:20 - sending preset to Star #1 R=15.3

-didn't seem to send, IIF shows "Active" in Cyan

-doing stop AO

UT 09:24 - pressing GO again after re-loading it worked

UT 09:27 - ROBS is nearly impossble to use

UT 09:32 - camera is railed in X, rotator angle = 373

-tried telsvc -90 and it failed

-need to change script - PA by 90

UT 09:36 - sending preset

UT 09:37 - pointing check

UT 09:40 - sending preset again

Rotator = 272 degrees

AO Info: rotator angle 272, pupil acquisition taking a long time for bin 3. Bad pupil centering bin3

the bad pupil centering was likely because we were centering bin3 on something much too faint.
el 77, provided mag 15.3, OCAM mag 16.9, bin 4, freq 390, average pupil intensity 15, 2.9 photons/subapp

UT 09:58 - optimize gain doesn't seem to be working

UT 10:16 managed to close loop and start taking data

UT 10:18 - DET X +40" absolute - taking skyframes

UT 10:28 - sky frames done, moving back to field

UT 10:29 - vibrations may be affecting things ~50mas residual vibrations on PSF. Going to try without Gopt

UT 10:38 - PSF does look a bit better

-results match the Strehl calculator predictor for this star's parameters

UT 10:48 - MGM tests on this star...

UT 11:06 - sending preset to Star #2 R=15.47, online tools say 16.3

Rotator = 403 degrees (PA=90) bin 3 hitting camera lens limit (sent magnitude 15.47)
Resending preset for bin 4 since OCAM mag is 16.6, bin 4 anyways, to investigate if the camera lens will rail at luci rotator 403 in bin 4 as well. it did not, had lots of range to center pupils.

UT 11:14 - sending some position angle but with Rmag =16.51

el 64, provided mag 16.51 (catalog 15.47), OCAM mag 16.6, bin 4, freq 432, average pupil intensity 15, 3.0 photons/subapp

-now have hit camera lmit in Y

-run gopt, 155 coutns in core

Rotator = 401 degrees

Pupils are on with pupil diff of <0.1 but Y camera lens position railed at 100 with Diff x0.00, y-0.02,

Diff x0.52, =-0.22 with lens railed, open loop

Loops opened...

Made new script rotated 90 degrees to PA=0

UT 11:27 - sending new script

Rotator = 310 degrees

el 61, provided mag 16.51 (catalog 15.47), OCAM mag 16.6, bin 4, freq 432, average pupil intensity 15, 3.0 photons/subapp

UT 12:01 - OCAM seems to have vanished, Fabio working on it

UT 12:12 - the pupil centering in bin 4 is pretty noisy, spending a lot more time not on target than on it

UT 12:26 - sending preset to AO 183 R=12.9 PA+180

-fieldstop moved a bit - this is my fault, i did not turn flex comp off on last moves (used a preset script but didn't turn flex off)

El 82. Given mag 12.9, OCAM 13.4, bin 2 freq 873, average pupil intensity 29, 6.5 photons/subapp

UT 12:38 - GEIRS GUI is non-responsive, does not show new imaegs. Daemon is still running and we can take imaegs

UT 12:41 - putting LUCI-1 in safe config.

I'm worried about the particle monitor. I see a lot of bumps right now that indicates we should be closed.

UT 12;43 - calling summit no one is answering

UT 12;44 - closing now

UT 12:55 - vent doors still open, closing now before cals
Calibrations:

UT 12:58 - L1 0192-0201 - J-band (lights off/lights on) - do not use

L1 0202 - ignore

-mask looks off set, will need to align

U 13:14 - alignment complete

L1 0207-0216 - J band (lights off/on)

L1 0217-0226 - FeII (lights off / on )

L1 0227-0236 - H2 (lights off / on )

UT 13:32 - Obs2 non-responsive on X2GO - this is not my internet connection at home

-switching to OBS4

UT 13:35 - L1 0237-0246 - Br Gamma (lights off / on)

UT 13:40 - L1 0247-0256 - H-band (lights off / on )

UT 13:46 - L1 0257-0266 - K band (lights off / on)

UT 13:51 - putting LUCI-1 into a safe config

Good night and good luck!

-- BarryRothberg - 01 Oct 2020
Topic revision: r1 - 01 Oct 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback